# Actor-Critic Model Predictive Control: Differentiable Optimization meets Reinforcement Learning

### v7

@@ V-C Is our approach robust to changes in the dynamics?
V-D What do we gain from Model-Predictive Value Expansion?
In this section, we conduct empirical experiments to evaluate the effect of the AC-MPC extended with Model-Predictive Value Expansion algorithm (introduced in Section III-H) in terms of training performance. To isolate the effect of MPVE and enable clear reward evaluation, we employ a straightforward hovering task for the drone. Here, the drone starts at a random position and orientation within a 1x1x1 meter cube around a designated equilibrium point (0, 0, 5 meters). The objective is to stabilize the drone at this point with zero velocity and a perfectly upright orientation (hover). We compare the training performance of AC-MPC-MPVE against AC-MLP. Both algorithms are trained with a fixed sample budget of 50 steps, corresponding to roughly 2 million samples collected from the environment. Because the MPVE algorithm primarily affects the learning of the critic, which uses the Generalized Advantage Estimation (GAE) algorithm [125], we explore the impact of our extension for various values of the temporal difference parameter $\lambda$ in TD($\lambda$). The parameter $\lambda$ determines the trade-off between bias and variance when estimating returns. A lower value of $\lambda$ favors lower variance but higher bias by relying more on immediate rewards, while higher values reduce bias but increase variance by incorporating more long-term information. To investigate this, we experiment with six different values for $\lambda$: 0.0, 0.2, 0.4, 0.6, 0.8, and 1.0. For each value, we repeat the training process three times with different random seeds.
Our results (visualized in Fig. 7) demonstrate that the sample efficiency gains achieved by MPVE are more significant for lower lambda values. This can be explained by the way the value function leverages data from the rollout buffer. Higher lambda values incorporate data from a larger portion of the rollout, increasing the influence of Monte Carlo estimates and reducing the relative impact of the information provided by MPVE predictions. In contrast, lower lambda values lead to stronger benefits from MPVE. This suggests that our algorithm is particularly advantageous for scenarios where the value function is trained with TD(0). This approach is often preferred when memory limitations exist (as it avoids storing the entire rollout buffer) or for non-episodic tasks.
@@ V-D What do we gain from Model-Predictive Value Expansion?
Figure 7: Reward at 50 steps of training for AC-MLP and AC-MPC (MPVE), depending on the $\lambda$ used in the TD($\lambda$). The policies are trained with a sample budget of 50 steps and the final reward is recorded, for different values of $\lambda$, for both AC-MPC (MPVE) and AC-MLP.
Figure 8: Reward evolution for the cases of $\lambda=0.0$ and $\lambda=0.6$, for AC-MLP and for AC-MPV (MPVE).
V-E Can we interpret the internal workings of AC-MPC?
This section investigates a potential connection between the critic network’s output, $V(\bm{x})$, and the learned cost matrix $Q(\bm{x})$. To simplify the analysis, we introduce three key assumptions. First, we assume that the observations $s_{k}$ are equivalent to the state of the system $x_{k}$. Second, the reward function is assumed to be quadratic in the state, expressed as $r(\bm{x})=-\bm{x}^{T}\bm{Q}\bm{x}$. This formulation captures the objective of stabilizing the system around a hover point. And third, we restrict the Model Predictive Control (MPC) horizon length to $T=2$. Under these simplifications, if we had a linear system with no constraints, the problem becomes equivalent to a Linear Quadratic Regulator (LQR) problem. In LQR, the value function is the solution to the Riccati equation, which is inherently quadratic with respect to the state.
@@ V-E Can we interpret the internal workings of AC-MPC?
Figure 9: Empirical probing of the correlation between the Hessian and the gradient of the learned critic $V(x)$ and the learned differentiable MPC cost terms $Q_{N}$ and $p_{N}$. (A): We train a linear classifier to infer $Q_{N}$ and $p_{N}$ from the Hessian of $V(x)$ and show the prediction from different 2D projections of $Q_{N}$ and $p_{N}$. We show how the ground truth can be predicted with high accuracy by only using a small linear classifier, which indicates that the data learned by the critic and $Q_{N}$ and $p_{N}$ are highly correlated. (B): Top: 3D plot of $\tau_{N}^{T}Q_{N}\tau_{N}+\tau_{N}^{T}p_{N}$ for the state dimensions of $v_{x}$ and $v_{y}$ after training. Bottom: 3D plot of $V(x)$ for the same dimensions. Apart from the change in sign and in scale, the similarity between both shapes indicates a high correlation between the value function and the learned cost parameters.
@@
where $\begin{bmatrix}A_{N-1},B_{N-1}\end{bmatrix}$ are the matrices associated to the linearized dynamics at timestep $N-1$, and $x_{N}=f(\tau_{N-1})$, $\frac{\partial V(x_{N})}{\partial x_{N}}$ is the gradient of the value function with respect to its inputs. For $H_{V}$:
@@
[…]
@@
In order to empirically show the relationship between the Hessian of the value function and what is learned in $Q_{N}$ and $p_{N}$, we perform a similarity study using linear probes. Particularly, we first train an AC-MPC agent to perform the stabilization at hover task, introduced in Section V-D. We choose this task here because, since the reward function is quadratic, it results in an optimization landscape that is easier to interpret. Since we need the derivative of the value function with respect to the full state, the critic needs to be trained with full state information. After training, we collect a dataset of uniformly distributed $5242882^{19}=$ 524288 datapoints, where every datapoint corresponds to a sample of $Q_{N}$, $p_{N}$, $H_{V}$ and $\Delta_{V}$, as per Eq. (V-E) and (V-E). Then, with this dataset, we train a single linear layer (without non-linearity) to predict the elements of the learned $Q_{N}$ and $p_{N}$ matrices from the elements of the $H_{V}$ and $p_{V}$, respectively. In the hypothesized case where the information contained in the learned cost is related to the information encoded by the gradient and Hessian of the value function, we expect the regression to have a high accuracy.
@@
V-F How sensitive is our approach to exploration hyperparameters?
Since our approach incorporates a model of the dynamics within the optimization problem, it has access to prior information before the training phase, and hence leverages prior knowledge of the dynamics in its architecture. This raises questions about the extent of exploration needed to discover an optimal policy and how the exploration parameter can be minimized to achieve optimal performance. To explore this, we study the effect of the initial policy standard deviation hyperparameter on the performance of both AC-MLP and AC-MPC. This hyperparameter influences the exploration behavior of the RL algorithms. In Fig. 10, we present reward plots for the training on the Vertical track under different values of this hyperparameter. The results show that AC-MLP requires a higher degree of exploration to identify the optimal solution. This indicates that AC-MPC effectively utilizes the prior dynamics knowledge, resulting in improved performance with reduced exploration. More interestingly, Fig. 10 shows that AC-MPC is less sensitive to the choice of this hyperparameter and yields high performance despite very low exploration values. On the other hand, AC-MLP fails to reach high-reward regions when trained with low values and generally underperforms in comparison to the proposed method.
@@ V-F How sensitive is our approach to exploration hyperparameters?
V-G How does our method perform when deployed in the real-world?
@@
Figure 11: Top view of the performance of AC-MLP and AC-MPC policies deployed in both nominal and a realistic simulator for the SplitS track. The policies are tested for 64 different initial positions, distributed in a cube of $0.5m$ of side length.
@@
Additionally, we perform a lap time comparison in the SplitS track. We train both AC-MPC and AC-MLP in the task of drone racing and execute different runs in two different simulation fidelities: nominal dynamics and realistic simulator. This realistic simulator, called NeuroBEM, and based in blade element momentum theory, comes from [120]. In Fig. 11 we show a top view of the deployment of both AC-MLP and AC-MPC in both of the aforementioned simulators. The policies are tested for 64 different initial positions, contained in a cube of $0.5$ m of side length. Additionally, Fig. 12 illustrates trajectories flown in two different tracks (Circle track and SplitS track) in simulation and the real world. Not only can our policy transfer to the real world without any fine-tuning, but it can do so while at the limit of handling, achieving speeds up to 21 m/s.
In Table II, we show the results of this comparison in terms of lap time and success rate, including both simulation modalities and real-world results.
TABLE II: Performance comparison between AC-MLP and AC-MPC in the SplitS track
@@
As observed in the lap time results, both AC-MPC and AC-MLP exhibit similar performance in terms of lap time, indicating that they are largely on par in this aspect. This similarity is expected, given that their asymptotic performance and final reward values are quite close, as illustrated in Figure 4. However, it is important to note that our approach not only achieves on-par performance but also demonstrates additional robustness to out-of-distribution scenarios (see Fig. 5, in Horizontal and Vertical tracks), and changes in the system dynamics (see Fig. 6, in SplitS track). This indicates that our method maintains its effectiveness under a wider range of conditions and uncertainties, while still matching the performance of the neural network-based architecture.
V-H How do training and inference times compare to standard actor-critic architectures?
In Table III, we show the training times (SplitS track) and the forward pass times for AC-MLP and for the proposed AC-MPC for different horizon lengths. As can be seen, while AC-MPC offers performance advantages, it incurs a higher computational cost. Training time is increased by a factor of approximately 30 (for a horizon length of N=2) due to the batched optimization required for each forward and backward pass. This computational overhead also affects online performance, resulting in a slower forward pass. Although the current Python implementation [^1] https://locuslab.github.io/mpc.pytorch/ of the differentiable MPC algorithm [26] achieves a sufficient deployment frequency of 50Hz for our system, optimized C/C++ implementations would significantly reduce training time and enable higher frequency deployments. Note, however, that this computational overhead did not affect closed-loop performance, as the 13.5ms execution time remained well within our 20ms control time step budget necessary to deploy the system successfully in the real world.
TABLE III: Solve times and inference times for different variations.
@@ VI Discussion and Conclusion
This work presented a new learning-based control framework that combines the advantage of differentiable model predictive control with actor-critic training. Furthermore, we showed that AC-MPC can leverage the prior knowledge embedded in the system dynamics to achieve better training performance (sample efficiency, robustness to hyperparameters) than the AC-MLP baseline, and to better cope with out-of-distribution scenarios, and with variations in the nominal dynamics, without any further re-training. Through empirical analysis, the interpretable nature of the differentiable MPC embedded in the actor-critic architecture allows us to conduct and empirical analysis that revealed a relationship between the learned value function and the learned cost functions. This provides a deeper understanding of the interplay between the reinforcement learning (RL) and MPC, offering valuable insights for future research. Additionally, our approach achieved zero-shot sim-to-real transfer, demonstrated by successfully controlling a quadrotor at velocities of up to 21 m/s in the physical world, being on par in performance with the AC-MLP baseline. In general, we show that our method can tackle challenging control tasks and achieves robust control performance for agile flight.
However, there are some limitations to be mentioned and to be improved in the future. First, an analytic model of the system is required for the differentiable MPC block, which limits our approach to mainly systems where the dynamics are known beforehand. Furthermore, training AC-MPC takes significantly longer than AC-MLP (see Table III), due to the fact that an optimization problem needs to be solved for both the forward and the backward pass through the actor network. In fact, there are open-source libraries [99, 100] that are recently evolving and implementing more efficient versions of differentiable MPC. Showcasing the approach in different tasks with different robots is another future direction.
@@
We believe that our method represents an important step in the direction of generalizability and robustness in RL. The proposed method demonstrates that modular solutions, which combine the advantages of learning-centric and model-based approaches, are becoming increasingly promising. Our approach potentially paves the way for the development of more robust RL-based systems, contributing positively towards the broader goal of advancing AI for real-world robotics applications.
Acknowledgments
We would like to thank Brandon Amos, for sharing his insights regarding the differentiable MPC code, and Sebastien Gros, for taking the time to proofread our manuscript. We would also like to thank Jiaxu Xing, Leonard Bauersfeld and Alex Barden for the insightful discussions and support with the hardware setup.

### v8

@@ V-D Is our approach robust to changes in the dynamics?
V-E What do we gain from Model-Predictive Value Expansion?
In this section, we conduct empirical experiments to evaluate the effect of the AC-MPC extended with Model-Predictive Value Expansion algorithm (introduced in Section III-H) in terms of training performance. To isolate the effect of MPVE and enable clear reward evaluation, we employ a straightforward hovering task for the drone. Here, the drone starts at a random position and orientation within a 1x1x1 meter cube around a designated equilibrium point (0, 0, 5 meters). The objective is to stabilize the drone at this point with zero velocity and a perfectly upright orientation (hover). We compare the training performance of AC-MPC-MPVE against AC-MLP and AC-MPC. All algorithms are trained with a fixed sample budget of 50 steps, corresponding to roughly 1.25 million samples collected from the environment. For these experiments, the horizon of the differentiable MPC is set to $N=3$, and the MPVE $H$ is set to the same value. Because the MPVE algorithm primarily affects the learning of the critic, which uses the Generalized Advantage Estimation (GAE) algorithm [130], we explore the impact of our extension for various values of the temporal difference parameter $\lambda$ in TD($\lambda$). The parameter $\lambda$ determines the trade-off between using the Bellman error and using the Monte-Carlo estimates for the value function approximation. A lower value of $\lambda$ favors lower variance but higher bias by relying more on immediate rewards, while higher values reduce bias but increase variance by incorporating more long-term information. To investigate this, we experiment with six different values for $\lambda$: 0.0, 0.2, 0.4, 0.6, 0.8, and 1.0. For each value, we repeat the training process three times with different random seeds.
Our results (visualized in Fig. 8) demonstrate that the sample efficiency gains achieved by MPVE are more significant for lower lambda values. This can be explained by the way the value function leverages data from the rollout buffer. Higher lambda values incorporate data from a larger portion of the rollout, increasing the influence of Monte Carlo estimates and diluting the relative impact of the information provided by MPVE predictions. In contrast, lower lambda values lead to stronger benefits from MPVE. This suggests that our algorithm is particularly advantageous for scenarios where the value function is trained with TD(0). This approach is often preferred when memory limitations exist (as it avoids storing the entire rollout buffer) or for non-episodic tasks.
@@ V-E What do we gain from Model-Predictive Value Expansion?
We would like to note that while we modify the value of $\lambda$ in this section for the sake of exposing the influence of MPVE, the value of $\lambda$ in the rest of the article is kept constant and equal to $\lambda=0.95$. This is a very common choice of $\lambda$ in general TD($\lambda$) architectures and empirically gives the most stable learning when roll-outs are long and the task demands a large planning horizon, as is the case in drone racing. As one can see in Fig. 8, as $\lambda$ approaches one, the returns of all three controllers become indistinguishable: with an almost Monte-Carlo target, the critic already receives a near-complete estimate of the tail cost, so the extra predictions injected by MPVE provide little additional benefit. Table II shows an overview of the hyperparameters used for training both AC-MPC and AC-MLP for the rest of the article.
TABLE II: AC-MLP and AC-MPC training hyper-parameters
Setting Value
Discount factor $\gamma$ 0.98
GAE $\lambda$ 0.95
Steps per update 250
Mini-batch size 25 000
SGD epochs 10
Clip range 0.2
Learning rate $3{\times}10^{-4}\!\rightarrow\!10^{-5}$
Entropy coef. 0.001
Value-loss coef. 0.5
Grad. clip 0.5
Init. log-std (AC-MPC / AC-MLP) 1.2-1.2 / 0.5-0.5
Policy net size [512, 512]
Value net size [512, 512]
Additionally, in table III we added the training time difference between adding MPVE and not adding it. It is important to note that since the actor network is not modified, the inference time remains identical for both algorithms.
TABLE III: Training time of the two algorithms for 2M steps.
Method Time
AC-MPC-MPVE 1 h 19 min
AC-MPC 0 h 50 min
Figure 8: For the quadrotor stabilization task, we show the reward at 50 steps of training for AC-MLP, AC-MPC and AC-MPC (MPVE), depending on the $\lambda$ used in the TD($\lambda$). The policies are trained with a sample budget of 50 steps and the final reward is recorded, for different values of $\lambda$, for both AC-MPC (MPVE) and AC-MLP.
Figure 9: Reward evolution for the quadrotor stabilization task, for the cases of $\lambda=0.0$ and $\lambda=0.6$, for AC-MLP and for AC-MPC (MPVE).
V-F Can we interpret the internal workings of AC-MPC?
This section investigates a potential connection between the critic network’s output, $V(\bm{x})$, and the learned cost matrix $Q(\bm{x})$. To simplify the analysis, we introduce three key assumptions. First, we assume that the observations $s_{k}$ are equivalent to the state of the system $x_{k}$. Second, the reward function is assumed to be quadratic in the state, expressed as $r(\bm{x})=-\bm{x}^{T}\bm{Q}\bm{x}$. This formulation captures the objective of stabilizing the system around a hover point. And third, we restrict the Model Predictive Control (MPC) horizon length to $T=2$. These assumptions are chosen to create a simplified, tractable setting that mirrors the structure of a classic Linear Quadratic Regulator (LQR) problem, where the relationship between the value function and the quadratic cost is well-understood. While our system remains non-linear, this setup allows us to isolate and probe the fundamental connection between the critic’s learned value function and the actor’s learned MPC cost.
@@ V-F Can we interpret the internal workings of AC-MPC?
Figure 10: Empirical probing of the correlation between the Hessian and the gradient of the learned critic $V(x)$ and the learned differentiable MPC cost terms $Q_{N}$ and $p_{N}$, for the task of quadrotor stabilization. (A): We train a linear classifier to infer $Q_{N}$ and $p_{N}$ from the Hessian of $V(x)$ and show the prediction from different 2D projections of $Q_{N}$ and $p_{N}$. We show how the ground truth can be predicted with high accuracy by only using a small linear classifier, which indicates that the data learned by the critic and $Q_{N}$ and $p_{N}$ are highly correlated. (B): Top: 3D plot of $\tau_{N}^{T}Q_{N}\tau_{N}+\tau_{N}^{T}p_{N}$ for the state dimensions of $v_{x}$ and $v_{y}$ after training. Bottom: 3D plot of $V(x)$ for the same dimensions. Apart from the change in sign and in scale, the similarity between both shapes indicates a high correlation between the value function and the learned cost parameters.
In this formulation, the term $V^{*}(x_{N})$ represents the optimal cost-to-go from the end of the finite horizon N. The hypothesis of our analysis is that the critic, through reinforcement learning, learns an approximation of this infinite-horizon value function. The actor, in turn, learns the quadratic cost parameters ($Q_{N},p_{N}$) that best represent the local, second-order approximation of this value function at the horizon’s boundary.
@@
where $\begin{bmatrix}A_{N-1},B_{N-1}\end{bmatrix}$ are the matrices associated to the linearized dynamics at timestep N1N-1, and $x_{N}=f(\tau_{N-1})$, $\frac{\partial V(x_{N})}{\partial x_{N}}$ is the gradient of the value function with respect to its inputs. For $H_{V}$:
@@
$+\frac{\partial V(x_{N})}{\partial x_{N}}\cancelto{0}{\frac{\partial^{2}f(\tau_{N-1})}{\partial\tau_{N-1}^{2}}}$
@@
In order to empirically show the relationship between the Hessian of the value function and what is learned in $Q_{N}$ and $p_{N}$, we perform a similarity study using linear probes. Particularly, we first train an AC-MPC agent to perform the stabilization at hover task, introduced in Section V-E. We choose this task here because, since the reward function is quadratic, it results in an optimization landscape that is easier to interpret. Since we need the derivative of the value function with respect to the full state, the critic needs to be trained with full state information. After training, we collect a dataset of uniformly distributed $2^{19}=524288$ datapoints, where every datapoint corresponds to a sample of $Q_{N}$, $p_{N}$, $H_{V}$ and $\Delta_{V}$, as per Eq. (V-F) and (V-F). Then, with this dataset, we train a single linear layer (without non-linearity) to predict the elements of the learned $Q_{N}$ and $p_{N}$ matrices from the elements of the $H_{V}$ and $p_{V}$, respectively. In the hypothesized case where the information contained in the learned cost is related to the information encoded by the gradient and Hessian of the value function, we expect the regression to have a high accuracy.
@@
V-G How sensitive is our approach to exploration hyperparameters?
Since our approach incorporates a model of the dynamics within the optimization problem, our policy has access to prior information before the training phase, and hence leverages prior knowledge of the dynamics in its architecture. In contrast, its AC-MLP counterpart consists of a randomly initialized MLP policy, which has no prior built-in knowledge about the system. This raises questions about the extent of exploration needed to discover an optimal policy and how the exploration parameter can be minimized to achieve optimal performance. To explore this, we study the effect of the initial policy standard deviation hyperparameter on the performance of both AC-MLP and AC-MPC. This parameter controls how much Gaussian noise is injected at the control input level of the policy. This means that for the AC-MPC, during training the noise is injected on top of the solution that comes from the optimizer. Therefore, excessive noise can result in these solver solutions being almost entirely disregarded. In Fig. 11, we present reward plots for the training on the Vertical track under different values of this hyperparameter. The results show that AC-MLP requires a higher degree of exploration to identify the optimal solution. This indicates that AC-MPC effectively utilizes the prior dynamics knowledge, resulting in improved performance with reduced exploration. More interestingly, Fig. 11 shows that AC-MPC is less sensitive to the choice of this hyperparameter and yields high performance despite very low exploration values. On the other hand, AC-MLP fails to reach high-reward regions when trained with low values and generally underperforms in comparison to the proposed method. As a conclusion of this study, the main guideline to tune the proposed AC-MPC is to use less exploration noise than generally needed for AC-MLP. Table II shows the hyperparameters used for both AC-MPC and AC-MLP, where the main difference is the initial exploration hyperparameter.
@@ V-G How sensitive is our approach to exploration hyperparameters?
V-H How does our method perform when deployed in the real-world?
@@
Figure 12: Top view of the performance of AC-MLP and AC-MPC policies deployed in both nominal and a realistic simulator for quadrotor racing in the SplitS track. The policies are tested for 64 different initial positions, distributed in a cube of $0.5m$ of side length.
@@
Additionally, we perform a lap time comparison in the SplitS track. We train both AC-MPC and AC-MLP in the task of drone racing and execute different runs in two different simulation fidelities: nominal dynamics and realistic simulator. This realistic simulator, called NeuroBEM, and based in blade element momentum theory, comes from [124]. We would like to note that quadcopters are complex robots and many difficulties arise when transferring policies from simulation to the real hardware, especially given the inherent aggressiveness of time-optimal policies. To address these limitations, we follow the sim2real techniques used in [24], where we first train policies in a non-fine-tuned simulator, then deploy these policies and collect more data. This data is then used to refine and develop a higher fidelity simulator, which is then used to train the final policy. On top of this, we use a significant amount of domain randomization in key parameters such as thrust map, mass, latency, drag and initial state. Finally, a term in the reward function that smoothes the control actions is needed to avoid jitter of the platform and ensure a safe deployment. In Fig. 12 we show a top view of the deployment of both AC-MLP and AC-MPC in both of the aforementioned simulators. The policies are tested for 64 different initial positions, contained in a cube of $0.5$ m of side length. Additionally, Fig. 13 illustrates trajectories flown in two different tracks (Circle track and SplitS track) in simulation and the real world. Not only can our policy transfer to the real world without any fine-tuning, but it can do so while at the limit of handling, achieving speeds up to 21 m/s.
In Table IV, we show the results of this comparison in terms of lap time and success rate, including both simulation modalities and real-world results.
TABLE IV: Performance comparison between AC-MLP and AC-MPC in the SplitS track
@@
As observed in the lap time results, both AC-MPC and AC-MLP exhibit similar performance in terms of lap time, indicating that they are largely on par in this aspect. This similarity is expected, given that their asymptotic performance and final reward values are quite close, as illustrated in Figure 5. However, it is important to note that our approach not only achieves on-par performance but also demonstrates additional properties in out-of-distribution scenarios (see Fig. 6, in Horizontal and Vertical tracks), and changes in the system dynamics (see Fig. 7, in SplitS track). This indicates that our method maintains its effectiveness under a wider range of conditions and uncertainties, while still matching the performance of the neural network-based architecture.
V-I How do training and inference times compare to standard actor-critic architectures?
In Table V, we show the training times (SplitS track) and the forward pass times for AC-MLP and for the proposed AC-MPC for different horizon lengths. As can be seen, while AC-MPC offers performance advantages, it incurs a higher computational cost. Training time is increased by a factor of approximately 30 (for a horizon length of N=2) due to the batched optimization required for each forward and backward pass. This computational overhead also affects online performance, resulting in a slower forward pass. Although the current Python implementation [^1] https://locuslab.github.io/mpc.pytorch/ of the differentiable MPC algorithm [31] achieves a sufficient deployment frequency of 50Hz for our system, optimized C/C++ implementations would significantly reduce training time and enable higher frequency deployments. Note, however, that this computational overhead did not affect closed-loop performance, as the 13.5ms execution time remained well within our 20ms control time step budget necessary to deploy the system successfully in the real world. Furthermore, although classical MPC often requires horizons longer than $N=2$, our architecture uses a trained critic network to learn and provide the long-term reasoning to the short-horizon solver. Our analysis in Section V-F shows that the learned MPC cost matrices are highly correlated with the Hessian of the critic’s value function. This indicates that the critic learns an accurate, infinite-horizon cost-to-go, which it then uses to inform the differentiable MPC.
TABLE V: Solve times and inference times for different variations.
@@ VI Discussion and Conclusion
This work presented a new learning-based control framework that combines the advantages of differentiable model predictive control with actor-critic training. Furthermore, we showed that for quadrotor racing tasks, AC-MPC can leverage the prior knowledge embedded in the system dynamics to achieve better training performance when compared to the main PPO baseline (AC-MLP), to better cope with out-of-distribution scenarios, and its ability to deal with variations in the nominal dynamics, without any further re-training. Through empirical analysis, the interpretable nature of the differentiable MPC embedded in the actor-critic architecture allows us to reveal a relationship between the learned value function and the learned cost functions. This provides a deeper understanding of the interplay between the reinforcement learning (RL) and MPC, offering valuable insights for future research. Additionally, our approach achieved zero-shot sim-to-real transfer, demonstrated by successfully controlling a quadrotor at velocities of up to 21 m/s in the physical world, being on par in performance with the AC-MLP baseline. In general, we show that our method can tackle challenging control tasks and achieves robust control performance for agile flight.
However, there are some limitations to be mentioned and to be improved in the future. First, AC‑MPC relies on differentiable MPC [31], whose theoretical framework and gradient derivation cover input constrained problems only. Therefore, state constraints are not supported, since this solver doesn’t implement them. Consequently, our controller inherits this limitation and can currently enforce input limits but not explicit state bounds. This becomes particularly evident in safety-related tasks, such as collision avoidance or enforcing a speed limit. In such scenarios, simply penalizing violations as a term in the reward function does not guarantee that the constraints will be respected at deployment time. We believe that the work introduced in this paper highlights and motivates the practical value of developing a differentiable MPC solver that natively supports state constraints Our results serve as strong encouragement for the numerical optimization community to further work in that direction. In addition, AC-MPC may underperform or fail in certain practical scenarios. One such case is solver non-convergence. Since AC-MPC relies on solving an optimization problem at every control step, it is susceptible to failures if the problem becomes ill-conditioned or infeasible due to, e.g., extreme cost weights, poor initialization, or tight constraints. Such failures can result in invalid control actions and incorrect gradient updates during training. While this issue was rare in our experiments (especially after enforcing positive definiteness of the cost matrices) it remains a relevant concern.
Another key limitation is the scalability of the optimization block. The differentiable MPC [31] solver used in this work is based on iterative LQR, which scales cubically with the state and input dimensions, and linearly with the horizon length. As a result, control of high-dimensional, more complex systems or use of longer horizons may become impractical for real-time deployment. While all experiments in this work are conducted using a quadrotor platform with moderate dimensionality, training and deploying and extending our AC-MPC framework to more complex, high-dimensional robotic systems represents a promising direction for future research. Such extensions would also motivate the development of more efficient differentiable solvers to further reduce training and inference times.
Furthermore, AC-MPC assumes that the system dynamics can be expressed as an analytical formulation, and that they are differentiable, since gradients must be propagated through the MPC solver during training. This restricts the applicability of our method to systems with known, differentiable dynamics expressed in closed form. In the absence of such models, the core assumptions needed to compute gradients through the solver no longer apply.
We believe that successfully tackling these challenges would unlock the demonstrated benefits of AC-MPC for a much wider range of real-world robotic applications, significantly broadening the impact of this architecture.
A final point of discussion is the distinction between AC-MPC and standard MBRL methods. First, standard MBRL is designed for scenarios where dynamics are unknown and must be learned from data, whereas the AC-MPC framework assumes access to a known, analytical dynamics model as a key component of its differentiable MPC module. Second, most state-of-the-art MBRL methods are off-policy, leveraging a replay buffer for improved sample efficiency. A third key difference lies in reward design. AC-MPC benefits from the flexibility to learn from any high-level reward signal, irrespective of its shape or differentiability. In contrast, MBRL methods that rely on planning through the model, particularly those using gradient-based techniques, often require more carefully structured or differentiable reward functions to ensure stability and enable effective planning.
In our work we show that our approach enhances an on-policy algorithm (PPO) by integrating an optimization-based prior from MPC. Given this, our approach is best understood as a method to improve the properties of an on-policy algorithm, and the most sensible comparisons are against the components it seeks to integrate: the pure model-free approach (AC-MLP) and the classic optimization-based controller (MPC). This allows for a clear demonstration of the value added by our hybrid architecture.
@@
While our results focus on agile quadrotor flight, we believe that the proposed method represents an important step in the direction of generalizability and out of distribution behaviour in RL. The proposed method demonstrates that modular solutions, which combine the advantages of learning-centric and model-based approaches, are becoming increasingly promising. Our approach potentially paves the way for the development of more robust RL-based systems, contributing positively towards the broader goal of advancing AI for real-world robotics applications.